# VisTR: Visualizations as Representations for Time-series Table Reasoning

### v2

@@ 4 Multimodal Visualization Alignment
We set out to establish a MLLM that aligns visualization modalities of chart, sketch, and text in a single joint embedding space. This section presents the data preparation, including data augmentation and user labeling strategies, to address the issue of lacking training dataset (Sect. 4.1), followed by modality binding for establishing a joint embedding space that binds chart, sketch, and text modalities together (Sect. 4.2). In the end, we present the implementation details (Sect. 4.3).
@@
4.1 Data Preparation
To train the multimodal visualization alignment model, we need to first prepare a dataset containing user intentions, which is not covered by current open-sourced datasets. To achieve the alignment of sketches, charts, and text, we employ two strategies to construct the training data: (1) data augmentation for chart-text pairing, and (2) user labeling for chart-sketch pairing. Figure 2 uses an example to illustrate the two data preparation strategies.
@@ 4.1 Data Preparation
4.2 Modality Binding
We aim to bind all text $\mathbb{T}$, chart $\mathbb{C}$ and sketch $\mathbb{S}$ modalities together in a joint embedding space. Existing CLIP-based works predominantly focus on aligning pairs of modalities, while our objective is to accomplish the alignment of three modalities in the absence of the pairings between text and sketch modalities. As shown in Figure 3, we employ a two-stage strategy: first, we employ modality-specific encoders to transform each modality into embeddings of uniform length[59]. All three modalities are processed using the Transformer architecture, leveraging its self-attention mechanism to effectively capture complex dependencies within each modality.
@@ 4.2 Modality Binding
$L_{2}(C,S)=[\alpha+H_{2}(C^{\prime},S)-H_{2}(C,S)]_{+}+[\alpha+H_{2}(C,S^{\prime})-H_{2}(C,S)]_{+},$
@@ 4.3 Implementation Details
5 VisTR System
With the MLLM for multimodal visualization alignment, we build VisTR system that allows users to intuitively explore, comprehend, and interact with tabular data using visualizations as the representation. This section details the implementation of visualization referencing (Sect. 5.1), visualization pruning (Sect. 5.2), and visualization interaction (Sect. 5.3) in VisTR system.
5.1 Visualization Referencing
@@
5.2 Visualization Pruning and Indexing
For a given table, we will generate a vast collection of visualization references, each of which is subsequently mapped into vector representations (i.e., embeddings). The large number of visualization references poses significant challenges to efficient storage and retrieval. To address these challenges, we design a combined approach of visualization pruning and indexing, as illustrated in Figure 4, to improve VisTR’s ability to manage and retrieve visualization references.
@@ 5.2 Visualization Pruning and Indexing
Visualization Pruning. Here, the key idea is to filter out less informative visualization references in advance while retaining those that provide significant insights. The main challenge lies in discerning and preserving highly valuable visualization references during pruning. For all visualization references for a variable $d_{i}$, it may encounter a scenario where two references exhibit similar change patterns, and one is entirely encompassed by the other. We quantify the similarity between visualization vectors using the Euclidean distance $d(c_{j},c_{k})$ between the two vectors $c_{j}$ and $c_{k}$. We introduce a pruning threshold to determine whether two visualization vectors are sufficiently similar to warrant pruning. If $d(c_{j},c_{k})$ is less than the set pruning threshold, the vector covering a shorter time range is filtered out. An example scenario is illustrated in Figure 4, where the ‘a peak’ change pattern exists in pink-boxed visualization references. In this case, the light-pink-boxed one has less information and is consequently filtered out.
@@
The pruning threshold acts as a predefined parameter that affects the number of stored visualization references. To assess its impact, we designed an experiment with four values: 0.5, 0.8, 1, and 1.5, and applied the thresholds to tables in a range of row numbers. Specifically, the row numbers were considered based on the typical sizes used in existing LLM-based table reasoning methods, which typically support data tables with fewer than 50 rows. For instance, Tabfact[7], a commonly used dataset containing over 16k Wiki tables, has an average of 12.96 rows per table (max.=48, min.=1, std.=8.46). Given the focus on time-series data tables and long-term trend analysis, we extended the experiment to tables with up to 1000 rows, approximately encompassing four trading years. We calculated the average number of visualization references across three sub-tables with corresponding table sizes extracted from the DJI. data table for each pruning threshold.
Figure 5 shows the relationship between the number of stored visualization references, the number of data table rows, and the pruning thresholds. When the pruning threshold is set to 1, the number of stored visualization references approximated 20% of the raw reference number. This aligns with the well-known Pareto principle (80/20 rule)[45], which is widely leveraged in various optimization strategies to balance efficiency and information retention. As such, we also set the pruning threshold to be 1. This strikes a balance between maintaining a rich representation set and reducing the storage footprint.
@@ 5.3 Visualization Interaction
Here, we employ two LLMs with in-context learning using prompts to decompose user queries and complete outputs respectively. Below we illustrate an example using the case shown in Figure VisTR: Visualizations as Representations for Time-series Table Reasoning(A). Given the query “What is the price trend of Apple during March?”, we first decompose the question into a retrieval task “return the visualization reference of Apple price in March”, and a cross-modal alignment task “recognize the trend in the retrieved visualization reference”. The corresponding visualization reference is first retrieved from Chroma, and then aligned with the ‘two peaks’ pattern to generate the final textual response “There is a two-peak price trend in Apple during March.”. Detailed prompts are provided in our open-sourced code.
Visual Interface. In addition to textual responses, VisTR also incorporates a user-friendly interface to enable intuitive table reasoning and exploration. As shown in Figure 6, the interface includes four main views: (A) Data Table, (B) Main View, (C) Chat Box, and (D) Pattern View. When a user uploads a table, the back-end of VisTR processes the tabular data into visualization references and stores them in the vector database, while the front-end interface displays the uploaded table in Data Table, and presents the visualization for the data in the Main View. By default, Main View uses an area graph for the overall context information and a line chart for the details. Various operations, such as selecting different variables or combinations of variables to display, are enabled in the drop-down box in the upper right corner of Main View. For the display of multi-variables, we use a multi-line chart with colors to encode different variables. The bottom area chart also allows for fixed-scale zooming and brushing, allowing users to observe detailed variable changes.
@@
6 Evaluation
To thoroughly evaluate VisTR, we conduct both quantitative evaluations and case studies. In the quantitative evaluations (Sect. 6.1), we assess the alignment ability of our fine-tuned MLLM. Subsequently, we present two cases on actual application scenarios (Sects. 6.2 & 6.3), representing a possible pathway through multimodal interactions to complete a time-series table reasoning task.
6.1 Effectiveness of Multimodal Visualization Alignment
@@ 6.1.1 Baselines and Metrics
6.1.2 Text-Chart Retrieval
@@
Table 1: Experimental results on chart-to-text retrieval test for all and individual chart types.
@@
Results. The quantitative results are presented in Table 1. Our fine-tuned MLLM demonstrates robust performance in aligning chart and text modalities, with overall accuracy and weighted F1-score exceeding 85% across all chart types. The results for each individual chart type also provide substantial evidence of the model’s effectiveness across various chart types. However, we observed that bar charts yield slightly lower metrics than other chart types. This discrepancy may be attributed to the inherent characteristics of bar charts, which utilize separated bars to encode data, resulting in less distinctiveness compared to the continuous line marks used in line charts and the filled area marks used in area charts. To enhance the accuracy of bar charts, factors such as bar widths may need to be considered during fine-tuning process.
@@
6.1.3 Sketch-Chart Retrieval
@@
As shown in Figure 7, our model achieved the highest similarity agreement among participants (Mean = 4.33, SD = 0.67), and the Unichart achieved the lowest agreement with Mean = 2.25 and SD = 1.17. Given the non-normal distribution of collected ratings, we initiated our analysis by conducting the Friedman test to evaluate the statistical significance of ratings among the four models. The p-value of the Friedman test is less than 0.01, indicating significant differences between models. To further pinpoint the specific differences among models, we employed the Wilcoxon signed-rank test for paired comparisons, accompanied by Bonferroni correction to control errors. The pairwise comparisons showed that our model significantly outperformed the others. The detailed comparison results are shown in Figure 7. The poor performance of UniChart may be due to its focus on OCR features rather than chart features, leading to a weaker grasp of the varied patterns present within charts and hand-drawn sketches. The comparison results underscore the importance of encoding patterns in our alignment model, confirm the judicious choice of CLIP-ViT as our base pre-trained model, and highlight the superiority of our fine-tuned model in binding sketch and chart modalities.
6.2 Case Study 1: Financial Data Exploration
@@
As shown in Figure 8, Mike explores the data as follows.
@@
6.3 Case Study 2: Air Pollutant Analysis
@@
Figure 9 illustrates Emily’s analysis process using VisTR to explore the multivariate data.
@@ 7 Discussion
  • General user queries. VisTR focuses on tasks related to reasoning about data change patterns in tables. Therefore, we primarily provide text-based user queries related to data change patterns, such as “what type of change pattern exists in the data?” However, this narrow focus also limits our ability to handle other general user queries, including those related to causal reasoning, such as “why did Apple stock rise in March?” When posed with such questions, VisTR may encounter the ‘hallucination’ issue, leading to incorrect queries and responses. To achieve a comprehensive exploration and reasoning of a data table, we require more general user queries, which may necessitate the incorporation of external knowledge sources.
@@
  • Intent-enhanced sketches. Users’ intuition about sketching targets often emerges from the exploration process. As a reason, VisTR offers insights into common patterns within the uploaded table. However, expressing complex and less common patterns solely through hand-drawn sketches can be challenging for users[32]. To address this challenge, several solutions can be considered. First, incorporating additional interactive features such as dragging, resizing, and erasing, sketches can provide users with greater flexibility in adjusting their intent[36]. Summarizing multi-attribute temporal patterns and introducing a more effective time-series segmentation method have shown promise in addressing this issue[51]. By implementing these methods, we can further enhance the capabilities of VisTR to support sketches with enhanced intents.
Scaling up the dataset. The successful alignment of our model heavily relies on acquiring high-quality training data. Scaling up the dataset to enhance the generalizability of VisTR presents significant challenges for the data augmentation and user labeling strategies currently employed. Collecting user-labeled sketches is time-consuming, and incorporating intents into sketches would further increase complexity. To address this limitation, one solution is to leverage crowdsourcing strategies, although this still requires significant manual effort. Another innovative technique would be to use large models to simulate and augment a variety of hand-drawn patterns. LLM-powered data augmentation has been a promising direction[11], and we are interested in using this approach for scaling up the dataset.
@@
8 Conclusion
This work introduces VisTR system, built upon a novel framework for table reasoning, expanding upon existing LLM-based table reasoning research. By leveraging visualizations as representations, the framework offers several advantages, including the ability to recognize data change patterns, mitigate LLM drift, and support visual-based exploration. The implementation of VisTR requires an effective multimodal visualization alignment model that binds chart, text, and sketch modalities together in a joint embedding space. This is accomplished by fine-tuning an MLLM with a new dataset of chart-text and chart-sketch pairings. Combined with strategies employed for visualization referencing, indexing, and pruning, we have developed an interactive system that facilitates multimodal interactions with time-series tables. The feasibility of VisTR is demonstrated through quantitative evaluations of the visualization alignment model and practical case studies on time-series data tables. This proposed framework offers a new research direction for table reasoning and enhances the fuzzy exploration process of existing LLM-based table reasoning methods. We aim to extend its support to a wider range of table types, accommodate more generalized user queries, and explore additional application scenarios in the future.
Acknowledgements.

### v3

@@ IV Multimodal Visualization Alignment
We set out to establish a MLLM that aligns visualization modalities of chart, sketch, and text in a single joint embedding space. This section presents the data preparation, including data augmentation and user labeling strategies, to address the issue of lacking training dataset (Sect. IV-A), followed by modality binding for establishing a joint embedding space that binds chart, sketch, and text modalities together (Sect. IV-B). In the end, we present the implementation details (Sect. IV-C).
@@
IV-A Data Preparation
To train the multimodal visualization alignment model, we need to first prepare a dataset containing user intentions, which is not covered by current open-sourced datasets. To achieve the alignment of sketches, charts, and text, we employ two strategies to construct the training data: (1) data augmentation for chart-text pairing, and (2) user labeling for chart-sketch pairing. Fig. 3 uses an example to illustrate the two data preparation strategies.
@@ IV-A Data Preparation
IV-B Modality Binding
We aim to bind all text $\mathbb{T}$, chart $\mathbb{C}$ and sketch $\mathbb{S}$ modalities together in a joint embedding space. Existing CLIP-based works predominantly focus on aligning pairs of modalities, while our objective is to accomplish the alignment of three modalities in the absence of the pairings between text and sketch modalities. As shown in Fig. 4, we employ a two-stage strategy: first, we employ modality-specific encoders to transform each modality into embeddings of uniform length[63]. All three modalities are processed using the Transformer architecture, leveraging its self-attention mechanism to effectively capture complex dependencies within each modality.
@@ IV-B Modality Binding
$L_{2}(C,S)=[\alpha+H_{2}(C^{\prime},S)-H_{2}(C,S)]_{+}$
$+[\alpha+H_{2}(C,S^{\prime})-H_{2}(C,S)]_{+},$ (3)
@@ IV-C Implementation Details
V VisTR System
With the MLLM for multimodal visualization alignment, we build VisTR system that allows users to intuitively explore, comprehend, and interact with tabular data using visualizations as the representation. This section details the implementation of visualization referencing (Sect. V-A), visualization pruning (Sect. V-B), and visualization interaction (Sect. V-C) in VisTR system.
V-A Visualization Referencing
@@
V-B Visualization Pruning and Indexing
For a given table, we will generate a vast collection of visualization references, each of which is subsequently mapped into vector representations (i.e., embeddings). The large number of visualization references poses significant challenges to efficient storage and retrieval. To address these challenges, we design a combined approach of visualization pruning and indexing, as illustrated in Fig. 5, to improve VisTR’s ability to manage and retrieve visualization references.
@@ V-B Visualization Pruning and Indexing
Visualization Pruning. Here, the key idea is to filter out less informative visualization references in advance while retaining those that provide significant insights. The main challenge lies in discerning and preserving highly valuable visualization references during pruning. For all visualization references for a variable $d_{i}$, it may encounter a scenario where two references exhibit similar change patterns, and one is entirely encompassed by the other. We quantify the similarity between visualization vectors using the Euclidean distance $d(c_{j},c_{k})$ between the two vectors $c_{j}$ and $c_{k}$. We introduce a pruning threshold to determine whether two visualization vectors are sufficiently similar to warrant pruning. If $d(c_{j},c_{k})$ is less than the set pruning threshold, the vector covering a shorter time range is filtered out. An example scenario is illustrated in Fig. 5, where the ‘a peak’ change pattern exists in pink-boxed visualization references. In this case, the light-pink-boxed one has less information and is consequently filtered out.
@@
The pruning threshold acts as a predefined parameter that affects the number of stored visualization references. To assess its impact, we designed an experiment with four values: 0.5, 0.8, 1, and 1.5, and applied the thresholds to tables of varying row counts. Specifically, the row numbers were considered based on the typical sizes used in existing LLM-based table reasoning methods, which typically support data tables with fewer than 50 rows. For instance, Tabfact[68], a commonly used dataset containing over 16k Wiki tables, has an average of 12.96 rows per table (max.=48, min.=1, std.=8.46). Given the focus on time-series data tables and long-term trend analysis, we extended the experiment to tables with up to 1000 rows, approximately encompassing four trading years. We calculated the average number of visualization references across three sub-tables with corresponding table sizes extracted from the DJI. data table for each pruning threshold.
Fig. 6 shows the relationship between the number of stored visualization references, the number of data table rows, and the pruning thresholds. When the pruning threshold is set to 1, the number of stored visualization references approximated 20% of the raw reference number. This aligns with the well-known Pareto principle (80/20 rule)[69], which is widely leveraged in various optimization strategies to balance efficiency and information retention. As such, we also set the pruning threshold to be 1. This strikes a balance between maintaining a rich representation set and reducing the storage footprint.
@@ V-C Visualization Interaction
Here, we employ two LLMs with in-context learning using prompts to decompose user queries and complete outputs respectively. Below we illustrate an example using the case shown in Fig. 1(A). Given the query “What is the price trend of Apple during March?”, we first decompose the question into a retrieval task “return the visualization reference of Apple price in March”, and a cross-modal alignment task “recognize the trend in the retrieved visualization reference”. The corresponding visualization reference is first retrieved from Chroma, and then aligned with the ‘two peaks’ pattern to generate the final textual response “There is a two-peak price trend in Apple during March.”. Detailed prompts are provided in our open-sourced code.
Visual Interface. In addition to textual responses, VisTR also incorporates a user-friendly interface to enable intuitive table reasoning and exploration. As shown in Fig. 7, the interface includes four main views: (A) Data Table, (B) Main View, (C) Chat Box, and (D) Pattern View. When a user uploads a table, the back-end of VisTR processes the tabular data into visualization references and stores them in the vector database, while the front-end interface displays the uploaded table in Data Table, and presents the visualization for the data in the Main View. By default, Main View uses an area graph for the overall context information and a line chart for the details. Various operations, such as selecting different variables or combinations of variables to display, are enabled in the drop-down box in the upper right corner of Main View. For the display of multi-variables, we use a multi-line chart with colors to encode different variables. The bottom area chart also allows for fixed-scale zooming and brushing, allowing users to observe detailed variable changes.
@@
VI Evaluation
To thoroughly evaluate VisTR, we conduct both quantitative evaluations and case studies. In the quantitative evaluations (Sect. VI-A), we assess the alignment ability of our fine-tuned MLLM. Subsequently, we present two classic cases on actual application scenarios (Sects. VI-B & VI-C), representing a possible pathway through multimodal interactions to complete a time-series table reasoning task.
VI-A Effectiveness of Multimodal Visualization Alignment
@@ VI-A1 Baselines and Metrics
VI-A2 Text-Chart Retrieval
@@
TABLE I: Experimental results on chart-to-text retrieval test for all and individual chart types.
@@
Results. The quantitative results are presented in Table. I. Our fine-tuned MLLM demonstrates robust performance in aligning chart and text modalities, with overall accuracy and weighted F1-score exceeding 85% across all chart types. The results for each individual chart type also provide substantial evidence of the model’s effectiveness across various chart types. However, we observed that bar charts yield slightly lower metrics than other chart types. This discrepancy may be attributed to the inherent characteristics of bar charts, which utilize separated bars to encode data, resulting in less distinctiveness compared to the continuous line marks used in line charts and the filled area marks used in area charts. To enhance the accuracy of bar charts, factors such as bar widths may need to be considered during fine-tuning process.
@@
VI-A3 Sketch-Chart Retrieval
@@
As shown in Fig. 8, our model achieved the highest similarity agreement among participants (Mean = 4.33, SD = 0.67), and the Unichart achieved the lowest agreement with Mean = 2.25 and SD = 1.17. Given the non-normal distribution of collected ratings, we initiated our analysis by conducting the Friedman test to evaluate the statistical significance of ratings among the four models. The p-value of the Friedman test is less than 0.01, indicating significant differences between models. To further pinpoint the specific differences among models, we employed the Wilcoxon signed-rank test for paired comparisons, accompanied by Bonferroni correction to control errors. The pairwise comparisons showed that our model significantly outperformed the others, as detailed in Fig. 8. The poor performance of UniChart may be due to its focus on OCR features rather than chart features, leading to a weaker grasp of the varied patterns present within charts and hand-drawn sketches. The comparison results underscore the importance of encoding patterns in our alignment model, confirm the judicious choice of CLIP-ViT as our base pre-trained model, and highlight the superiority of our fine-tuned model in binding sketch and chart modalities.
VI-B Case Study 1: Financial Data Exploration
@@
As shown in Fig. 9, Mike explores the data as follows.
@@
VI-C Case Study 2: Air Pollutant Analysis
@@
Fig. 10 illustrates Emily’s analysis process using VisTR to explore the multivariate data.
@@ VII Discussion
  • General queries. VisTR focuses on tasks related to reasoning about data change patterns in tables. Therefore, we primarily provide text-based user queries related to data change patterns, such as “what type of change pattern exists in the data?” However, this narrow focus also limits our ability to handle other general user queries, including those related to causal reasoning, such as “why did Apple stock rise in March?” When posed with such questions, VisTR may encounter the ‘hallucination’ issue, leading to incorrect queries and responses. To achieve a comprehensive exploration and reasoning of a data table, we require more general user queries, which may necessitate the incorporation of external knowledge sources.
  • Dateset bias. The successful alignment of our model heavily relies on acquiring high-quality training data. However, the dataset used for fine-tuning, while demonstrating its effectiveness, is limited in size and diversity, posting challenges to generalizability. This limitation becomes particularly evident when handling complex or unconventional sketch representations, where a richer and more diverse dataset is essential for robust model fine-tuning. The dataset collection process itself introduces potential biases and inconsistencies. Scaling up the dataset to address these limitations presents significant challenges. First, the limited number of users contributing to user-labeled sketches may embed subjective preferences into the chart-sketch-text pairings. While crowd-sourcing strategies offer a way to accelerate data collection, they are prone to similar subjective biases and the quality control for such large-scale manual efforts remains a significant challenge. An innovative alternative lies in data augmentation techniques to reduce reliance on manual sketch collection. Second, the filtered trend words from textual descriptions may oversimplify visual information, neglecting nuances in semantics. To enrich textual descriptions, a potential solution proposed by Time-LLM[76] is to generate text prototypes by assembling candidate vocabularies. Then the generated rich text descriptions can be paired with charts through crowd-sourced rating frameworks[77].
@@
  • Intent-enhanced sketches. Users’ intuition about sketching targets often emerges from the exploration process. As a reason, VisTR offers insights into common patterns within the uploaded table. However, expressing complex or uncommon patterns solely through hand-drawn sketches can be challenging for users[22]. To address this challenge, several solutions can be considered. First, incorporating additional interactive features like dragging, resizing, and erasing can provide users with greater flexibility in adjusting their intent[78]. Summarizing multi-attribute temporal patterns and introducing a more effective time-series segmentation method have shown promise in addressing this issue[79]. By implementing these methods, we can further enhance the capabilities of VisTR to support sketches with enhanced intents.
VisTR vs. Traditional Methods. VisTR introduces an innovative attempt to convert time-series tables into visual representations, enabling efficient multi-modal binding and continuous context reasoning. Compared to traditional subsequence or sketch matching methods, such as metric-based[41], regular expression-based[74] or deep-learning-based approaches[72], VisTR processes varying data facet lengths into uniform vision tokens, offering smoother multi-modal interaction and significant resource savings due to reduced prompt lengths[80]. However, visual representations come with notable limitations. They often underperform traditional metric-based methods in matching precision, incur higher storage costs, and may ignore shorter time subsequences in visualization referencing module. Additionally, fine-tuning models for aligning modalities of chart, sketch, and text is resource-intensive. Despite these challenges, the rapid advancements of multi-modal large language models (MLLMs) make this research direction promising. Future efforts should focus on improving precision and optimizing storage efficiency to fully leverage the potential of visual representations in advancing multi-modal frameworks.
@@
VIII Conclusion
We present VisTR, a new framework for time-series table reasoning that leverages visualizations as representations to bridge raw data and human cognition. By enabling enhanced pattern recognition, mitigating context drift, and supporting visual-based exploration, VisTR addresses key limitations of existing LLM-based methods. VisTR integrates a fine-tuned MLLM, trained on a newly constructed dataset of chart-text and chart-sketch pairs, to align charts, text, and sketches in a unified embedding space. Combined with strategies for visualization referencing, pruning, and indexing, VisTR provides an interactive system for seamless multimodal interactions with time-series tables. Quantitative evaluations and case studies demonstrate its effectiveness in aligning multimodal inputs and improving reasoning accuracy. VisTR opens new research directions for integrating visualizations into table reasoning and enhancing multimodal data exploration. Future work will extend its applicability to broader table types, generalized queries, and diverse application scenarios.